OpenStudy (anonymous):

Identify all true statements √9 is a rational # √-1 is NOT a rational # √ab=√a *√b √a+b=√a+√b 9y√3 and 9√3y are like radicals ^3√5 is the symbol for the square root of 5 √a-b ≠ √a-√b

terenzreignz (terenzreignz):

Miss Butterfly... we meet again ;) Shall we go through them one by one?

OpenStudy (anonymous):

hiiiiii yes please i need to get this one right to pass my assigment please help

terenzreignz (terenzreignz):

Let's start with \[\large \sqrt9 \ \text{ is a rational number }\] What exactly is a rational number?

terenzreignz (terenzreignz):

A rational number is a number that you can express as a fraction, with both its numerator and denominator integers. So, can you do that with the square root of 9?

OpenStudy (anonymous):

yes by 3?

terenzreignz (terenzreignz):

First of all, what exactly is the square root of 9?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

heeeeelp

terenzreignz (terenzreignz):

Sorry. Was preoccupied. 3... which is an integer, right?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Are not all integers also rational numbers?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Identify all true statements √9 is a rational # √-1 is NOT a rational # √ab=√a *√b √a+b=√a+√b are these all true statements?

terenzreignz (terenzreignz):

Two of them are not always true...